# Color Bind: Exploring Color Perception in Text-to-Image Models

### v2

@@ Abstract
Text-to-image generation has recently seen remarkable success, granting users with the ability to create high-quality images through the use of text. However, contemporary methods face challenges in capturing the precise semantics conveyed by complex multi-object prompts. Consequently, many works have sought to mitigate such semantic misalignments, typically via inference-time schemes that modify the attention layers of the denoising networks. However, prior work has mostly utilized coarse metrics, such as the cosine similarity between text and image CLIP embeddings, or human evaluations, which are challenging to conduct on a larger-scale. In this work, we perform a case study on colors—a fundamental attribute commonly associated with objects in text prompts, which offer a rich test bed for rigorous evaluation. Our analysis reveals that pretrained models struggle to generate images that faithfully reflect multiple color attributes—far more so than with single-color prompts—and that neither inference-time techniques nor existing editing methods reliably resolve these semantic misalignments. Accordingly, we introduce a dedicated image editing technique, mitigating the issue of multi-object semantic alignment for prompts containing multiple colors. We demonstrate that our approach significantly boosts performance over a wide range of metrics, considering images generated by various text-to-image diffusion-based techniques. We will make our code, benchmark and evaluation protocol publicly available.
@@ 1 Introduction
Hans Hofmann, perhaps one of the most influential painters of the 20th century, famously stated that: “In nature, light creates the color. In the picture, color creates the light.” According to Hofmann, our natural world can be described in its fullest expression through color. And indeed, colors—and particularly how they are perceived and described by humans—have been extensively studied in cognitive science [17, 4, 35]. The recent rise of foundational computational models has invited parallel studies [1, 24] that explore whether colors are encoded within these models similarly to the human brain. In this work, we are interested in exploring color perception in the context of modern text-to-image models. These models are capable of turning our imagination into stunning high-quality images, perhaps analagous to Hofmann’s “picture”, begging the question: Do these models perceive colors just like us?
Unfortunately, the answer might be no. Evidence suggests that these models struggle in faithfully generating the semantics conveyed by complex multi-subject prompts [9, 33, 10]. Consider the results depicted in Figure LABEL:fig:teaser. While text-to-image models mostly succeed in generating a “cornflower-blue gift box”, they are challenged by the more complicated prompt that also contains additional colored entities. We observe that a case study on colors offers a unique opportunity for performing a rigorous evaluation over text-to-image generative models. By contrast to coarse CLIP-based metrics, entangled VQA-based scores [20], or small-scale human evaluations conducted by prior work, performance can be measured explicitly using perceptual color differences.
@@
Accordingly, we introduce a training-free approach that utilizes attention-based diffusion models for editing existing images to match color specifications. In particular, we propose two objectives for guiding our optimization scheme: an attention loss for binding the colors to the right object and a color loss that forces objects to have the right color. A key insight enabling our proposed attention loss is that a simplified color-less reference text prompt can provide supervision for grounding noisy attention maps. Inspired by the psychological Stroop effect [25] which suggests that incongruent stimuli such as color words printed in differing colors (e.g. “blue” printed in red) have a stronger interference effect, our loss is built on the premise that misalignments in attribute binding leads to errors in the attention maps. Hence, we propose a loss that encourages the cross-attention maps of the full prompt to be more congruent to the simplified reference prompt, thus allowing for binding the colors to the correct image regions.
@@ 2 Related Work
Misalignment Mitigation in Text-to-image Models. Pretrained text-to-image diffusion models [36, 40, 8, 12] have garnered significant attention due to their remarkable ability to generate diverse high-quality images. However, a critical shortcoming of existing text-to-image models pertains to their limited capacity in representing the exact semantics, particularly in complicated multi-subject prompts.
@@
Prior work [9, 33] has analyzed and identified typical misalignments, such as catastrophic neglect (i.e., failure to generate one or more mentioned subjects) and incorrect attribute binding (i.e., failure to match between attributes and subjects). Various mitigation schemes have been proposed. These increase semantic alignment by modifying the text embeddings [13, 38] or by better aligning the cross-attention maps [9, 33]. Several techniques mitigating semantic misalignments also utilize additional inputs, such as fine-grained segmentation maps [19], coarse spatial layouts [32, 10] or through the use of compositional generation [23]. In this work, we focus on the problem of incorrect attribute binding, focusing specifically on colors. We introduce a new benchmark for studying such semantic misalignments, analyzing the performance of existing text-to-image models on our proposed benchmark.
@@
Color Perception in Computational Models. Understanding color perception and its relation with associated color names in language is a longstanding goal in cognitive science [17, 4, 35]. With the rise of foundational computational models, one line of work has explored the question of whether this alignment is also reflected in these models. In particular, prior work has analyzed color perception in text-only language models. Abdou et al. [1] propose probing techniques for measuring semantic alignment of basic color terms. Loyola et al. [24] extend this methodology to more abstract and complex color descriptions, such as “NYC Taxis”. In our work, we are interested in exploring color perception in multimodal text-to-image models, which are grounded in the visual world. Hence, we can measure this semantic alignment directly, without the need to compare high-dimensional language encodings with their corresponding CIELAB color values.
@@
Another line of work has explored color understanding for comparing multimodally-trained models with text-only models. Alper et al. [3] demonstrated that vision-and-language models can better associate objects with colors, in comparison to text only models, like BERT [11]. Grimal et al. [15] proposed a new metric for evaluating alignment in text-to-image models, highlighting issues of color leakage and significant performance drops when prompts involve two colors. Several work [28, 22] have studied the effect of reporting bias on language models, demonstrating that they are able to connect object and colors despite the reporting bias in text [26]. Monroe et al.[27] conducted a study on colors in context, training classifers to predict colors given multiple colors as reference, demonstrating that color term have a broader meaning when in context with other colors. Kim et al. [18] noted that diffusion models are less effective at modeling color-related attributes compared to shape-related attributes. Butt et al [6] also focus on colors in the context of text-to-image generative models. They propose an approach for optimizing tokens representing target colors. In this work, we study to what extent existing models can depict target colors, considering single and multiple objects and colors and propose an editing framework which does not require training per RGB color value.
@@
Text-guided Image Editing. The generative power of text-to-image diffusion models has also fueled increasing interests in text-guided image editing [16, 7, 29, 21, 41, 5, 42]. To edit input images, existing methods typically manipulate the internal representations of the denoising networks, in particular the self and cross attention maps [37, 2] or employs instruction-based editing. To mitigate color misalignments in text-to-image models, we propose an editing method that builds upon the attention-based framework presented in Liu et al. [21]. We adapt it to our problem setting, as further detailed in Section 4.
@@ 3 The CompColor Benchmark
We aim to better understand the limitations of current image generation models in responding to complex prompts with multiple fine-grained attributes. However, a key challenge here is the lack of quantitative metrics that can accurately capture such fine-grained semantic alignment. Prior work typically relies on coarse metrics, such as the cosine similarity between the text prompt and the generated image in CLIP embedding space (denoted as $\text{CLIP}_{\text{Sim}}$), or the visual question-answering capabilities of vision-and-language models (e.g., $\text{VQA}_{\text{Score}}$ [20]). As shown in Fig 2, these metrics often fail in multi-object, multi-color scenarios—scoring misaligned generations as highly relevant.
To directly evaluate fine-grained semantic alignment in generated images, we introduce a benchmark composed of complex, multi-object prompts with carefully specified color attributes, which we refer to as CompColor. We focus on color as a first diagnostic axis for compositionality: it is a commonly described visual property, closely tied to object identity, and uniquely suited for precise, perceptually grounded quantification—having been extensively studied in cognitive science and computer vision [31, 1, 24]. We can therefore use Euclidean distance in the CIELAB color space, a perceptually uniform metric, to explicitly quantify color-object alignment in generated images [1, 31, 6]. This allows for a systematic and scalable analysis of models’ compositional performance. We next describe our benchmark construction; additional details and visualizations are provided in the supplementary material.
@@
Object Selection for Paired Prompts. We follow prior works on attribute leakage [9, 14, 33] and take the union of objects used in Attend-and-Excite, Rich Text, and SynGen. From this union, we carefully remove: (i) Objects with dominant intrinsic colors (e.g., bananas are mostly yellow). (ii) Objects with small or irregular surfaces (e.g., glasses only have border color). For each color pair, we randomly select two objects and generate five composition prompts. We also conduct experiments with prompts containing more than two colored objects, included in the supplementary.
@@ 3.1 Evaluation Metrics
To evaluate the performance of the model in adhering to color specifications, we introduce an explicit color evaluation pipeline. First, we segment the target object using SAM [34]. To reduce noise and capture dominant colors, we quantize the colors on the segmented object by performing k-means clustering and find the cluster closest to the ground-truth color specified in the prompt.
@@
To supplement this with a more fine-grained evaluation, we also compute two additional metrics: (1) LAB L2 Distance, measuring Euclidean distance between the matched cluster color and the ground-truth color in the 3D CIELAB color space, a perceptually uniform color representation, similarly to [1]. (2) RGB L2 Distance Euclidean distance between the matched cluster color and the ground truth color in the RGB space, similarly to [14]. Extended results over these metrics are provided in the supplementary.
@@ 3.2 Baselines
    Stable Diffusion (SD) [36]: A widely used UNet-based text-to-image diffusion model. We use versions 1.4, 1.5, and 2.1 without further modifications.
@@
    FLUX [12] is a rectified flow based method that integrates a transformer backbone, aiming to capture global text–image context more effectively. We use FLUX-dev for evaluation.
@@
    Attend-and-Excite (A&E) [9], optimizes the latents by encouraging to maximize the cross-attention maps values for each subject token. It is based on SD 1.4.
@@
    Structred-Diffusion (StructDiff) [13], incorporates linguistic structures with the diffusion guidance process based on the controllable properties of manipulating cross-attention layers. It is based on SD 1.4.
@@
    SynGen [33] uses language-driven cross-attention losses to force the cross-attention map of the modifier to largely overlap with the cross-attention map of the noun, while remaining largely disjoint with the maps corresponding to other nouns and modifiers. It is based on SD 1.4.
@@
    RichText [14] enables precise color control of generated objects by enforcing a color loss on the intermediate latents masked by the object mask. It is based on SD 1.5.
@@
    Bounded-Attention (BA) [10] is a layout conditional generation model based on SD 1.5.
@@ 4.1 Preliminaries
We begin by providing brief background about text-to-image diffusion models, which will form the backbone of our approach. We use text-to-image diffusion models for generating an initial image. That is, given a prompt $P$, these diffusion models generate an image $I$ that matches the prompt. Modern models operate in the latent space of an autoencoder (typically a fixed VQ-VAE [36]): they generate a latent tensor $Z$ that can be decoded into $I$.
@@ 4.2.1 Binding colors to the right object
As previously discussed, modern generative models struggle with binding the colors to the right object. This corroborates past work that has demonstrated this binding problem for general attributes object [9, 33]. The reason for this difficulty is apparent from the cross-attention maps (see Figure 3): the color tokens (e.g., “pink”) attend to the wrong parts of the image (in this case the flowers instead of the spoon). Thus, intuitively we can ensure proper binding of colors/attributes to the right objects if we “correct” these cross-attention maps. But how do we make this correction?
We propose an approach that constructs pseudo-ground-truth for the cross attention maps, different from the contrastive loss proposed by prior work [33] which can degrade performance when objects are close or share similar attributes (as shown in Section 3). We take inspiration from Stroop probing [3], which is an approach for probing language representations. Stroop probing analyzes language models by comparing the representations of a reference prompt (e.g., ”a photo of a banana”) to a prompt with colors specified (e,g, ”a photo of a blue banana”). Here we take inspiration from this idea and propose to compare the cross-attention maps of our prompt with color (e.g., ”a photo of a blue banana”) with the cross-attention maps of a simplified prompt (” a photo of a banana”). Ideally, we would want the cross-attention maps of ”blue” and ”banana” from the former to match the cross-attention maps of ”banana” from the latter. Thus, the cross-attention maps from the simplified prompt provide pseudo-ground-truth cross-attention maps for all the objects and attributes.

### v3

@@ Abstract
Text-to-image generation has recently seen remarkable success, granting users with the ability to create high-quality images through the use of text. However, contemporary methods face challenges in capturing the precise semantics conveyed by complex multi-object prompts. Consequently, many works have sought to mitigate such semantic misalignments, typically via inference-time schemes that modify the attention layers of the denoising networks. However, prior work has mostly utilized coarse metrics, such as the cosine similarity between text and image CLIP embeddings, or human evaluations, which are challenging to conduct on a larger-scale. In this work, we perform a case study on colors—a fundamental attribute commonly associated with objects in text prompts, which offer a rich test bed for rigorous evaluation. Our analysis reveals that pretrained models struggle to generate images that faithfully reflect multiple color attributes—far more so than with single-color prompts—and that neither inference-time techniques nor existing editing methods reliably resolve these semantic misalignments. Accordingly, we introduce a dedicated image editing technique, mitigating the issue of multi-object semantic alignment for prompts containing multiple colors. We demonstrate that our approach significantly boosts performance over a wide range of metrics, considering images generated by various text-to-image diffusion-based techniques. Our code, benchmark and evaluation protocol is publicly available on our project webpage.
@@ 1 Introduction
Hans Hofmann, perhaps one of the most influential painters of the 20th century, famously stated that: “In nature, light creates the color. In the picture, color creates the light.” According to Hofmann, our natural world can be described in its fullest expression through color. And indeed, colors—and particularly how they are perceived and described by humans—have been extensively studied in cognitive science [kay1978linguistic, berlin1991basic, regier2007color]. The recent rise of foundational computational models has invited parallel studies [abdou2021languagemodelsencodeperceptual, loyola2023perceptualstructureabsencegrounding] that explore whether colors are encoded within these models similarly to the human brain. In this work, we are interested in exploring color perception in the context of modern text-to-image models. These models are capable of turning our imagination into stunning high-quality images, perhaps analagous to Hofmann’s “picture”, begging the question: Do these models perceive colors just like us?
Unfortunately, the answer might be no. Evidence suggests that these models struggle in faithfully generating the semantics conveyed by complex multi-subject prompts [chefer2023attendandexciteattentionbasedsemanticguidance, NEURIPS2023_0b08d733, dahary2024yourselfboundedattentionmultisubject]. Consider the results depicted in Figure LABEL:fig:teaser. While text-to-image models mostly succeed in generating a “cornflower-blue gift box”, they are challenged by the more complicated prompt that also contains additional colored entities. We observe that a case study on colors offers a unique opportunity for performing a rigorous evaluation over text-to-image generative models. By contrast to coarse CLIP-based metrics, entangled VQA-based scores [lin2024evaluatingtexttovisualgenerationimagetotext], or small-scale human evaluations conducted by prior work, performance can be measured explicitly using perceptual color differences.
@@
Accordingly, we introduce a training-free approach that utilizes attention-based diffusion models for editing existing images to match color specifications. In particular, we propose two objectives for guiding our optimization scheme: an attention loss for binding the colors to the right object and a color loss that forces objects to have the right color. A key insight enabling our proposed attention loss is that a simplified color-less reference text prompt can provide supervision for grounding noisy attention maps. Inspired by the psychological Stroop effect [macleod1991half] which suggests that incongruent stimuli such as color words printed in differing colors (e.g. “blue” printed in red) have a stronger interference effect, our loss is built on the premise that misalignments in attribute binding leads to errors in the attention maps. Hence, we propose a loss that encourages the cross-attention maps of the full prompt to be more congruent to the simplified reference prompt, thus allowing for binding the colors to the correct image regions.
@@ 2 Related Work
Misalignment Mitigation in Text-to-image Models. Pretrained text-to-image diffusion models [rombach2022highresolutionimagesynthesislatent, xue2024raphael, cao2024controllable, esser2024scalingrectifiedflowtransformers] have garnered significant attention due to their remarkable ability to generate diverse high-quality images. However, a critical shortcoming of existing text-to-image models pertains to their limited capacity in representing the exact semantics, particularly in complicated multi-subject prompts.
@@
Prior work [chefer2023attendandexciteattentionbasedsemanticguidance, NEURIPS2023_0b08d733] has analyzed and identified typical misalignments, such as catastrophic neglect (i.e., failure to generate one or more mentioned subjects) and incorrect attribute binding (i.e., failure to match between attributes and subjects). Various mitigation schemes have been proposed. These increase semantic alignment by modifying the text embeddings [feng2022training, tunanyan2023multi] or by better aligning the cross-attention maps [chefer2023attendandexciteattentionbasedsemanticguidance, NEURIPS2023_0b08d733]. Several techniques mitigating semantic misalignments also utilize additional inputs, such as fine-grained segmentation maps [kim2023dense], coarse spatial layouts [phung2024grounded, dahary2024yourselfboundedattentionmultisubject] or through the use of compositional generation [liu2023compositionalvisualgenerationcomposable]. In this work, we focus on the problem of incorrect attribute binding, focusing specifically on colors. We introduce a new benchmark for studying such semantic misalignments, analyzing the performance of existing text-to-image models on our proposed benchmark.
@@
Color Perception in Computational Models. Understanding color perception and its relation with associated color names in language is a longstanding goal in cognitive science [kay1978linguistic, berlin1991basic, regier2007color]. With the rise of foundational computational models, one line of work has explored the question of whether this alignment is also reflected in these models. In particular, prior work has analyzed color perception in text-only language models. Abdou et al. [abdou2021languagemodelsencodeperceptual] propose probing techniques for measuring semantic alignment of basic color terms. Loyola et al. [loyola2023perceptualstructureabsencegrounding] extend this methodology to more abstract and complex color descriptions, such as “NYC Taxis”. In our work, we are interested in exploring color perception in multimodal text-to-image models, which are grounded in the visual world. Hence, we can measure this semantic alignment directly, without the need to compare high-dimensional language encodings with their corresponding CIELAB color values.
@@
Another line of work has explored color understanding for comparing multimodally-trained models with text-only models. Alper et al. [alper2023bertblindexploringeffect] demonstrated that vision-and-language models can better associate objects with colors, in comparison to text only models, like BERT [devlin2018bert]. Grimal et al. [grimal2024tiammetricevaluating] proposed a new metric for evaluating alignment in text-to-image models, highlighting issues of color leakage and significant performance drops when prompts involve two colors. Several work [paik2021worldoctopusreportingbias, liu2022largeroctopiamplifyreporting] have studied the effect of reporting bias on language models, demonstrating that they are able to connect object and colors despite the reporting bias in text [misra2016seeinghumanreportingbias]. Monroe et al.[monroe2017colorscontextpragmaticneural] conducted a study on colors in context, training classifers to predict colors given multiple colors as reference, demonstrating that color term have a broader meaning when in context with other colors. Kim et al. [kim2024attributebasedinterpretableevaluation] noted that diffusion models are less effective at modeling color-related attributes compared to shape-related attributes. Butt et al [butt2024colorpeelcolorpromptlearning] also focus on colors in the context of text-to-image generative models. They propose an approach for optimizing tokens representing target colors. In this work, we study to what extent existing models can depict target colors, considering single and multiple objects and colors and propose an editing framework which does not require training per RGB color value.
@@
Text-guided Image Editing. The generative power of text-to-image diffusion models has also fueled increasing interests in text-guided image editing [hertz2022prompttopromptimageeditingcross, Cao_2023_ICCV, Patashnik_2023_ICCV, liu2024understandingcrossselfattentionstable, yu2025anyeditmasteringunifiedhighqualityz, brooks2023instructpix2pixlearningfollowimage, zhang2024magicbrushmanuallyannotateddataset]. To edit input images, existing methods typically manipulate the internal representations of the denoising networks, in particular the self and cross attention maps [tumanyan2023plug, alaluf2024cross] or employs instruction-based editing. To mitigate color misalignments in text-to-image models, we propose an editing method that builds upon the attention-based framework presented in Liu et al. [liu2024understandingcrossselfattentionstable]. We adapt it to our problem setting, as further detailed in Section 4.
@@ 3 The CompColor Benchmark
We aim to better understand the limitations of current image generation models in responding to complex prompts with multiple fine-grained attributes. However, a key challenge here is the lack of quantitative metrics that can accurately capture such fine-grained semantic alignment. Prior work typically relies on coarse metrics, such as the cosine similarity between the text prompt and the generated image in CLIP embedding space (denoted as $\text{CLIP}_{\text{Sim}}$), or the visual question-answering capabilities of vision-and-language models (e.g., $\text{VQA}_{\text{Score}}$ [lin2024evaluatingtexttovisualgenerationimagetotext]). As shown in Fig 2, these metrics often fail in multi-object, multi-color scenarios—scoring misaligned generations as highly relevant.
To directly evaluate fine-grained semantic alignment in generated images, we introduce a benchmark composed of complex, multi-object prompts with carefully specified color attributes, which we refer to as CompColor. We focus on color as a first diagnostic axis for compositionality: it is a commonly described visual property, closely tied to object identity, and uniquely suited for precise, perceptually grounded quantification—having been extensively studied in cognitive science and computer vision [10.1007/978-3-319-10602-1_7, abdou2021languagemodelsencodeperceptual, loyola2023perceptualstructureabsencegrounding]. We can therefore use Euclidean distance in the CIELAB color space, a perceptually uniform metric, to explicitly quantify color-object alignment in generated images [abdou2021languagemodelsencodeperceptual, 10.1007/978-3-319-10602-1_7, butt2024colorpeelcolorpromptlearning]. This allows for a systematic and scalable analysis of models’ compositional performance. We next describe our benchmark construction; additional details and visualizations are provided in the supplementary material.
@@
Object Selection for Paired Prompts. We follow prior works on attribute leakage [chefer2023attendandexciteattentionbasedsemanticguidance, Ge_2023_ICCV, NEURIPS2023_0b08d733] and take the union of objects used in Attend-and-Excite, Rich Text, and SynGen. From this union, we carefully remove: (i) Objects with dominant intrinsic colors (e.g., bananas are mostly yellow). (ii) Objects with small or irregular surfaces (e.g., glasses only have border color). For each color pair, we randomly select two objects and generate five composition prompts. We also conduct experiments with prompts containing more than two colored objects, included in the supplementary.
@@ 3.1 Evaluation Metrics
To evaluate the performance of the model in adhering to color specifications, we introduce an explicit color evaluation pipeline. First, we segment the target object using SAM [ravi2024sam2]. To reduce noise and capture dominant colors, we quantize the colors on the segmented object by performing k-means clustering and find the cluster closest to the ground-truth color specified in the prompt.
@@
To supplement this with a more fine-grained evaluation, we also compute two additional metrics: (1) LAB L2 Distance, measuring Euclidean distance between the matched cluster color and the ground-truth color in the 3D CIELAB color space, a perceptually uniform color representation, similarly to [abdou2021languagemodelsencodeperceptual]. (2) RGB L2 Distance Euclidean distance between the matched cluster color and the ground truth color in the RGB space, similarly to [Ge_2023_ICCV]. Extended results over these metrics are provided in the supplementary.
@@ 3.2 Baselines
    Stable Diffusion (SD) [rombach2022highresolutionimagesynthesislatent]: A widely used UNet-based text-to-image diffusion model. We use versions 1.4, 1.5, and 2.1 without further modifications.
@@
    FLUX [esser2024scalingrectifiedflowtransformers] is a rectified flow based method that integrates a transformer backbone, aiming to capture global text–image context more effectively. We use FLUX-dev for evaluation.
@@
    Attend-and-Excite (A&E) [chefer2023attendandexciteattentionbasedsemanticguidance], optimizes the latents by encouraging to maximize the cross-attention maps values for each subject token. It is based on SD 1.4.
@@
    Structred-Diffusion (StructDiff) [feng2022training], incorporates linguistic structures with the diffusion guidance process based on the controllable properties of manipulating cross-attention layers. It is based on SD 1.4.
@@
    SynGen [NEURIPS2023_0b08d733] uses language-driven cross-attention losses to force the cross-attention map of the modifier to largely overlap with the cross-attention map of the noun, while remaining largely disjoint with the maps corresponding to other nouns and modifiers. It is based on SD 1.4.
@@
    RichText [Ge_2023_ICCV] enables precise color control of generated objects by enforcing a color loss on the intermediate latents masked by the object mask. It is based on SD 1.5.
@@
    Bounded-Attention (BA) [dahary2024yourselfboundedattentionmultisubject] is a layout conditional generation model based on SD 1.5.
@@ 4.1 Preliminaries
We begin by providing brief background about text-to-image diffusion models, which will form the backbone of our approach. We use text-to-image diffusion models for generating an initial image. That is, given a prompt $P$, these diffusion models generate an image $I$ that matches the prompt. Modern models operate in the latent space of an autoencoder (typically a fixed VQ-VAE [rombach2022highresolutionimagesynthesislatent]): they generate a latent tensor $Z$ that can be decoded into $I$.
@@ 4.2.1 Binding colors to the right object
As previously discussed, modern generative models struggle with binding the colors to the right object. This corroborates past work that has demonstrated this binding problem for general attributes object [chefer2023attendandexciteattentionbasedsemanticguidance, NEURIPS2023_0b08d733]. The reason for this difficulty is apparent from the cross-attention maps (see Figure 3): the color tokens (e.g., “pink”) attend to the wrong parts of the image (in this case the flowers instead of the spoon). Thus, intuitively we can ensure proper binding of colors/attributes to the right objects if we “correct” these cross-attention maps. But how do we make this correction?
We propose an approach that constructs pseudo-ground-truth for the cross attention maps, different from the contrastive loss proposed by prior work [NEURIPS2023_0b08d733] which can degrade performance when objects are close or share similar attributes (as shown in Section 3). We take inspiration from Stroop probing [alper2023bertblindexploringeffect], which is an approach for probing language representations. Stroop probing analyzes language models by comparing the representations of a reference prompt (e.g., ”a photo of a banana”) to a prompt with colors specified (e,g, ”a photo of a blue banana”). Here we take inspiration from this idea and propose to compare the cross-attention maps of our prompt with color (e.g., ”a photo of a blue banana”) with the cross-attention maps of a simplified prompt (” a photo of a banana”). Ideally, we would want the cross-attention maps of ”blue” and ”banana” from the former to match the cross-attention maps of ”banana” from the latter. Thus, the cross-attention maps from the simplified prompt provide pseudo-ground-truth cross-attention maps for all the objects and attributes.